OpenStudy (anonymous):

Three-fourths a number decreased by one is between negative six and sixteen. Find all such numbers.

OpenStudy (anonymous):

\[-6<\frac{3}{4}x-1<16\]

OpenStudy (anonymous):

\[-5<\frac{3}{4}x<17\] \[-20<3x<68\] \[-\frac{20}{3}<x<\frac{68}{3}\]

OpenStudy (anonymous):

i guess "between" means maybe you should write \[-\frac{20}{3}\leq x \leq \frac{68}{3}\]

OpenStudy (anonymous):

hello my myininaya!

myininaya (myininaya):

hey

OpenStudy (anonymous):

did you see your personal thank you?

myininaya (myininaya):

yes i did

myininaya (myininaya):

im the best :)

OpenStudy (anonymous):

now that my relief is here i can get my puzzles and head off to bed. don't stay up all night!

myininaya (myininaya):

i just came for a small visit im going to play some games! :)

OpenStudy (anonymous):

yes \[\color{red}{\text{myininaya = best!}}\]

myininaya (myininaya):

you are the best too satellite!

OpenStudy (anonymous):

gamer girl! \[\color{\red}{\text{pumpin peaches apple pie}}\]